# ViewDelta: Text-Prompted Change Detection in Unaligned Images

### v1

@@ 4.3 Combating Inpainting Noise
Although inpainting produces seemingly realistic changes, we noticed that the inpainted regions tend to have inpainting artifacts (as observed in other works in the literature [6, 17, 7]). To discourage the model from simply learning these artifacts instead of the changes between images, we inpaint “red herring” masks. These are a randomly selected subset of tracked SAV masks that are classified as not the text prompt. These tracked masks are inpainted in Frame B but are not added to the change mask. During training, this forces the model to learn the change between the two images that are consistent with text prompt rather than labeling inpainting noise as changes.
@@ 5 Experiments
Figure 6: Qualitative Results: We show the segmentation predictions in purple of our model on the test sets, along with the ground truth in blue. All SYSU-CD test images use the text prompt of: “urban development, suburban expansion, pre-construction groundwork, vegetation alteration, road widening, and coastal construction”.
@@ 5.1 Quantitative Evaluation
ViewDelta 67.82 80.82 75.88 86.47
Table 1: Quantitative Results On The CSeg Dataset.
Model mIoU F1 Rec Prec
@@
ViewDelta 36.83 53.83 55.97 51.86
Table 2: Quantitative Results On The PSCD Dataset.
@@
FC-EF [8] 60.09 75.07 75.84 74.32
FC-Siam-diff [8] 56.96 72.57 61.21 89.13
FC-Siam-conc [8] 61.75 76.35 71.03 82.54
@@
ViewDelta 67.18 80.37 77.40 83.58
Table 3: Quantitative Results on SYSU-CD Dataset
We quantitatively evaluate the performance of ViewDelta on three diverse datasets: CSeg, PSCD [39], and SYSU-CD [40], comparing it against several state-of-the-art methods using standard metrics, including Intersection over Union (IoU), F1 score, Recall (Rec), and Precision (Prec). Results are summarized in Tabs. 1, 2 and 3. On the CSeg dataset Tab. 1, ViewDelta achieves an IoU of 67.82%, an F1 score of 80.82%, a recall of 75.88% with a high precision of 86.47%. For the PSCD dataset Tab. 2, which includes viewpoint changes and challenging illumination conditions, ViewDelta competes with CSCDNet+SSCDNet [39], achieving a mean IoU of 36.83%. We observed that the “object (traffic)” and “object (other)” labels in PSCD often had inconsistent or infrequent annotations, contributing to the relatively lower mIoU. Finally, on the SYSU-CD dataset (Tab. 3), which consists of satellite imagery, ViewDelta achieves state-of-the-art performance with an IoU of 67.18% and an F1 score of 80.37%, outperforming all compared methods on these metrics. These results demonstrate consistent improvements across datasets, underscoring the robustness of ViewDelta. We attribute this performance to the model’s multi-domain training, which enables it to generalize effectively across different change detection domains and scale more easily than specialized methods as combinations of datasets can be used to train with.
@@ 5.2 Qualitative Evaluation
We show our model results on example image pairs from the test sets of CSeg, PSCD, and SYSU-CD in Fig. 6. All SYSU-CD test images use the text prompt of: “urban development, suburban expansion, pre-construction groundwork, vegetation alteration, road widening, and coastal construction”.
@@ 5.3 Limitations
CSeg and PSCD are limited to minor perspective differences, which limits the models generalizability to large perspective changes. However, expanding the dataset to include larger perspective changes is challenging due to the limited availability of the required data in addition to the difficultly in labeling the data. Another limitation is the exponential increase in the number of image tokens required for higher resolution images that may contain small changes. However, this can be mitigated as compute is scaled.
@@ 6 Conclusion
We introduce ViewDelta, the first text-prompted change detection model that accepts unaligned images and a text prompt to output a binary segmentation of changes relevant to the user-provided text. We also release an accompanying dataset, CSeg, comprising 100,311 pairs of images with text prompts and corresponding change detection labels. We demonstrate the effectiveness of ViewDelta both quantitatively and qualitatively on datasets with a wide variety of images including indoor, outdoor, street level, synthetic, and satellite images.
Acknowledgments: Placeholder

### v2

@@ 4.2 Combating Inpainting Noise
Although inpainting produces seemingly realistic changes, we noticed that the inpainted regions tend to have inpainting artifacts (as observed in other works in the literature [9, 19, 10]). We use “red herring” masks to prevent the model from learning artifacts instead of image changes. These red herring masks are a randomly selected subset of zero to ten SA-1B masks that are not classified as any of the classes in the text prompt. Since SA-1B doesn’t provide mask classifications, we use Grounded SAM 2 with a low confidence threshold to ensure that the SA-1B mask doesn’t coincide with a change class instance. These red herring masks are inpainted on image B but not added to the change mask. During training, this forces the model to learn the change between the two images that are consistent with text prompt rather than labeling inpainting noise as changes. Adding red herring masks to “all” image pairs is not applicable, since every change needs to be accounted for which would include all inpaints.
4.3 Dataset Statistics
There is a total of 501,153 change image pairs in CSeg split into 451,090 pairs in train and 50,063 in test. There are 24,298 unique classes and 339,348 unique prompts in train and 7,285 unique classes and 7,326 unique prompts in test. There are 1,408 unique classes and 35,271 unique prompts not seen in train. 53,257 and 5,947 pairs were “all” image pairs in the train and test sets respectively at about a 12% of the total for each set. CSeg’s class counts are top heavy in that more common classes such as “clouds”, “sky”, “people”, “water”, “trees” and “roof” appear exponentially more often than less common classes like “calcite”, “aldravanda”, and “ice hammer” as depicted in Fig. 4. This is due to how the majority of images in SA-1B include these common objects. During class proposals, we select the classes that are least used in CSeg to mitigate this problem.
Although LLaVa-Next and Grounded SAM 2 are very impressive vision models, they cannot produce flawless class proposals and instance masks for all images. This consequently affects the quality of CSeg. To quantitatively measure the quality of CSeg, we manually check randomly sampled images until the percent accuracy clearly converges. We validated 500 images - about 0.1% of the entire dataset - and observed an accuracy of 94.0%. This results in a margin of error of ±2.22% at a 95% statistical confidence. The large majority of errors stem from Grounded SAM 2 yielding masks that are misclassified, especially with more challenging classes. We provide additional examples in Fig. 5 for further evaluation.
Figure 4: Class count distributions: top 1% most frequent classes and the complete dataset. These figures don’t include the “all” class. As class counts decrease, the class gets more impractical.
Figure 5: Evaluation of CSeg quality. Change mask label is highlighted in blue. Manual inspection of the generated images show a 94% agreement with a human reviewer.
@@ 5 Experiments
Figure 6: Qualitative Results: We show ViewDelta predictions in red on the test sets, along with the ground truth in blue. All SYSU-CD test images use the text prompt of: “urban development, suburban expansion, pre-construction groundwork, vegetation alteration, road widening, and coastal construction”.
@@ 5.1 Quantitative Evaluation
ViewDelta 85.91 92.42 89.25 95.82
Table 1: Quantitative Results On The CSeg Dataset by a jointly trained model, the checkpoint used is the same across all reported results for the general model. The CSeg test set contains prompts not seen in the training set.
Model IoU F1 Rec Prec
Fine-tuned models
@@
ViewDelta 55.54 78.31 69.47 73.47
General model
ViewDelta 52.24 68.63 72.57 65.10
Table 2: Quantitative Results on PSCD Dataset. The highest values are highlighted in bold and the general model results are from the same checkpoint across all results.
@@
Fine-tuned models
FC-EF [11] 61.04 75.81 75.17 76.47
FC-Siam-diff [11] 61.01 75.79 75.30 76.28
FC-Siam-conc [11] 60.23 75.18 76.75 73.67
@@
ChangeFormerV4 [4] 65.03 78.81 77.90 79.74
BIT-50 [6] 66.14 79.62 77.90 81.42
TransUNetCD [23] 66.79 80.09 77.73 82.59
SwinSUNet [49] 68.89 81.58 79.75 83.50
MambaBCD [7] 71.10 83.11 80.31 86.11
ViewDelta 70.09 82.41 79.91 85.08
General model
ViewDelta 68.59 81.37 75.42 88.33
Table 3: Quantitative Results on SYSU-CD Dataset. The highest values are highlighted for First and Second. The general model results are from the same checkpoint across all results. We additionally finetune the general ViewDelta model to evaluate its finetuning performance.
We quantitatively evaluate the performance of ViewDelta on three diverse datasets: CSeg, PSCD [36], and SYSU-CD [37], comparing it against several state-of-the-art methods using standard metrics, including Intersection over Union (IoU), F1 score, Recall (Rec), and Precision (Prec). Results are summarized in Tabs. 1, 2 and 3.
On the CSeg dataset Tab. 1, ViewDelta achieves excellent performance, with an IoU of 85.91%, an F1 score of 92.42%, a recall of 89.25%, and notably high precision of 95.82%.
For the PSCD dataset Tab. 2, which includes viewpoint changes and challenging illumination conditions, ViewDelta significantly outperforms prior fine-tuned methods, achieving an IoU of 55.54%, an F1 score of 78.31%, recall of 69.47%, and precision of 73.47%. The general model also shows strong performance with an IoU of 52.24%.
Finally, on the SYSU-CD dataset Tab. 3, consisting of satellite imagery, the ViewDelta fine-tuned model achieves competitive results with an IoU of 70.09% and an F1 score of 82.41%, closely following the state-of-the-art MambaBCD-Base method. The general model maintains strong performance, achieving an IoU of 68.59% and demonstrating superior precision at 88.33%.
These results demonstrate competitive performance across datasets, underscoring ViewDelta’s robustness. We attribute this success to the model’s multi-domain training approach. Given this strong performance from a relatively simple architecture, we anticipate that further architectural enhancements or specialized fine-tuning approaches could lead to even more substantial performance improvements in the future.
@@ 5.2 Qualitative Evaluation
We present qualitative results of our model on example image pairs from the CSeg, PSCD, and SYSU-CD test sets in Fig. 6. For the SYSU-CD test set, all images use the text prompt: “urban development, suburban expansion, pre-construction groundwork, vegetation alteration, road widening, and coastal construction.” For the PSCD test set, each image uses a prompt corresponding to the classes defined in the dataset, such as “vehicle” or “structure.”
5.3 Ablation Study
We perform extensive ablation studies to investigate the impact of various embedding strategies on the training efficiency of ViewDelta. First, we evaluate combinations of image and text embeddings on a random subset of the CSeg dataset containing 22,624 images, allowing models to train for 12 hours. Results are presented in Tab. 4.
Embeddings IoU
SigLip + Dinov2 56.44
SigLip + Patch Embedding 26.16
CLIP + Dinov2 53.46
CLIP + Patch Embedding 27.56
Table 4: Evaluation of embedding combinations on a subset of CSeg.
The results demonstrate that the combination of SigLip text embeddings and Dinov2 image embeddings consistently outperforms other configurations. Additionally, using Dinov2 embeddings significantly surpasses patch embeddings, regardless of text encoder choice.
Further, we evaluate the specific impact of frozen Dinov2 embeddings compared to fully-trained patch embeddings across multiple datasets simultaneously. Both configurations use SigLip for text embeddings, and results are summarized in Tab. 5.
Image Embedding CSeg PSCD SYSU-CD
Frozen Dinov2 85.91 52.24 68.59
Patch Embedding 61.05 16.87 34.81
Table 5: Impact of frozen Dinov2 versus trained patch embeddings across multiple datasets.
We observe that frozen Dinov2 embeddings significantly outperform fully-trained patch embeddings, indicating that pretrained image embeddings not only provide a more robust feature representation but also improve convergence speed and overall model generalization. This advantage is particularly pronounced given the dataset size limitations encountered in our training scenarios.
@@ 5.4 Limitations
CSeg is limited to synthetic perspective differences, which limits the models generalizability to large perspective changes. Expanding the dataset to include larger perspective changes is challenging due to the limited availability of the required data in addition to the difficultly in labeling the data. Lack of real data of image pairs with textual prompts. Another limitation is the exponential increase in the number of image tokens required for higher resolution images that may contain small changes. However, this can be mitigated as compute is scaled.
@@ 6 Conclusion
We presented ViewDelta, a novel multimodal architecture for text-prompted change detection capable of processing unaligned images through the integration of natural language prompts. By leveraging pretrained text and image embeddings, ViewDelta effectively generalizes across diverse domains without requiring retraining for novel changes. We introduced the CSeg dataset, a comprehensive dataset specifically crafted for text-conditioned change detection tasks, providing over 500,000 image pairs with associated textual prompts and annotated binary masks. Extensive experiments demonstrated that ViewDelta achieves robust and competitive performance across various challenging domains, including indoor, outdoor, street-level, synthetic, and satellite imagery. Our findings underscore the model’s flexibility and effectiveness, paving the way for future advancements in generalized, text-conditioned change detection methodologies.